For y = f(x) = 2x - 1, x = 5, and △x = 3 find 


a) △y for the given x and △x values, 

b) dy = f'(x)dx, 

c) dy for the given x and △x values.  
Answer #1

SOLUTION :


y = f(x) = 2x - 1


=> f’ (x) = 2   


a.


Now, 


f’ (x) = dy / dx ≈ ∆y / ∆x

=> ∆y ≈ f’ (x) ∆x 

=> ∆y (at x = 5, ∆x = 3) is  = 2 * 3

=> ∆y = 6  (ANSWER).


b.


dy = f’(x) dx = 2 dx (ANSWER).


c.


dy ≈ ∆y = 6 (ANSWER)

answered by: Tulsiram Garg